$SPAD/src/input char.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool char.output
)set message test on
)set message auto off
)clear all
-- Input generated from CharacterXmpPage
--S 1 of 13
chars := [char "a", char "A", char "X", char "8", char "+"]
--R
--R
--R   (1)  [a,A,X,8,+]
--R                                                        Type: List(Character)
--E 1
--S 2 of 13
space()
--R
--R
--R   (2)
--R                                                              Type: Character
--E 2
--S 3 of 13
quote()
--R
--R
--R   (3)  "
--R                                                              Type: Character
--E 3
--S 4 of 13
escape()
--R
--R
--R   (4)  _
--R                                                              Type: Character
--E 4

--S 5 of 13
[ord c for c in chars]
--R
--R
--R   (5)  [97,65,88,56,43]
--R                                                          Type: List(Integer)
--E 5
--S 6 of 13
[upperCase c for c in chars]
--R
--R
--R   (6)  [A,A,X,8,+]
--R                                                        Type: List(Character)
--E 6
--S 7 of 13
[lowerCase c for c in chars]
--R
--R
--R   (7)  [a,a,x,8,+]
--R                                                        Type: List(Character)
--E 7
--S 8 of 13
[alphabetic? c for c in chars]
--R
--R
--R   (8)  [true,true,true,false,false]
--R                                                          Type: List(Boolean)
--E 8
--S 9 of 13
[upperCase? c for c in chars]
--R
--R
--R   (9)  [false,true,true,false,false]
--R                                                          Type: List(Boolean)
--E 9
--S 10 of 13
[lowerCase? c for c in chars]
--R
--R
--R   (10)  [true,false,false,false,false]
--R                                                          Type: List(Boolean)
--E 10
--S 11 of 13
[digit? c for c in chars]

--R
--R
--R   (11)  [false,false,false,true,false]
--R                                                          Type: List(Boolean)
--E 11
--S 12 of 13
[hexDigit? c for c in chars]
--R
--R
--R   (12)  [true,true,false,true,false]
--R                                                          Type: List(Boolean)
--E 12
--S 13 of 13
[alphanumeric? c for c in chars]
--R
--R
--R   (13)  [true,true,true,true,false]
--R                                                          Type: List(Boolean)
--E 13
)spool
)lisp (bye)

———-

References^{[1] nothing}

6